Calculate the % sodium in a 0.721 g detergent sample. Final yield is  0.320 g NaP2O4 


1
Expert's answer
2021-02-09T03:46:23-0500

The mass of sodium in the NaP2O4 sample is:

m(Na) = Mr(Na) × m(NaP2O4) / Mr(NaP2O4) = 23 g/mol × 0.320 / 149 g/mol = 0.05 g

% Na = [m(Na) / m (detergent)] × 100% = [0.05 g / 0.721 g] × 100 % = 6.93 %


Answer: 6.93 %